MICROWAVE SPECTROSCOPY MEASUREMENTS OF THE GAS PHASE STRUCTURE OF CYCLOPEN-TADIENYL TUNGSTENTRICARBONYL HYDRIDE ^{a}

\underline{CHANDANA KARUNATILAKA},  CHAKREE  TANJAROON,  KRIS-

TEN S. KECK, AND STEPHEN G. KUKOLICH, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Microwave spectra for five unique ^{13}C isotopomers of cyclopentadienyltungstentricarbonyl hydride were measured using a Flygare-Balle type microwave spectrometer system. The new rotational constants were combined with the previously obtained rotational constants for normal and deuterium analogues for various tungsten isotopes. A total of fifty-seven rotational constants were acquired and used in the least squares fit to obtain the gas phase structure of this d^{4} - metal mono hydride complex. The results from the structural fit yielded the W-H bond length: r_{o}(W-H)= 1.788(20) Å, which agrees very well with the previously reported Kraitchman value of r_{s}(W-H)= 1.79(4) Å. The present study also yielded the distance from tungsten to the centroid distance of the C_{5}H_{5} ring: r(W-Cp)= 2.02(1) Å, the ring radius of Cp: r(Cp)= 1.2016(4) Å, the average bond length from tungsten to the carbonyl carbon, r(W-CO)= 1.981(14) Å, and the average cyclopentadienyl C-C bond length of 1.421(4) Å. Results obtained from the structural fit are in much closer agreement with the Kraitchman values than with the DFT results.

aSupported by THE NATIONAL SCIENCE FOUNDATION - CHE0304969